# Identity-related Speech Suppression in Generative AI Content Moderation

### v1

@@ 5.1 Automated content moderation APIs
Supp. Identity Supp. Identity
Google 1.21 non-chr. 3.50 non-chr.
Jigsaw 1.58 lgbt 1.20 non-chr.
OpenAI score 3.37 non-chr. 2.16 women
OpenAI flag 2.19 non-chr. 6.44 men
Llama Guard 2.33 white 1.79 non-chr.
Anthropic 1.97 white 3.10 men
@@
Table 4: Worst identity-related speech suppression values (Supp.) and associated identity group achieving that value (Identity) across content moderation APIs tested. OpenAI has speech suppression values associated with both scores (based on $i$-medians) and flags (based on $i$-FPRs) shown, while Google and Jigsaw are both score-based and Llama Guard and Anthropic are both flag-based.
Each instance of text in each dataset is sent through five automated content moderation APIs to receive its categorization and/or scored results. Jigsaw’s Perspective API is a traditional content moderation system used for moderation of online comment sections and other user-generated content. OpenAI [22] and Google [13] have automated content moderation APIs designed to be part of generative AI moderation, and Llama Guard (as provided by Octo AI [25]) and Anthropic [1] provide specific prompts (see Appendix Figures 2, 3, and 4) to give to generative AI systems to moderate user-generated content. Each system is designed for slightly different categorization aims and their outputs vary, ranging from toxicity to sexual content to politics (see Table 3).
The types of values returned by these moderation systems also vary, including overall flags indicating content deemed violating, per-category flags indicating content identified with that category, as well as overall or per-category scores (see Table 3). Some systems return more than one of these types of values. While Google’s returned per-category scores are meant to represent confidence that an input belongs to the associated category [13], and thus are comparable across categories, OpenAI’s per-category scores must be normalized to be comparable across categories [21]. We follow the normalization mechanism described in [21]; we determine score flagging thresholds (see Appendix Table 10) per category and divide by these thresholds so that scores above $1.0$ flag. While a previous audit of OpenAI determined that multiple runs per instance were useful for the stability of the results [21], we run each instance through each API only once given the size of the combined datasets and the cost of some APIs.
5.2 Speech Suppression results
For each of the five APIs, we collect content moderation scores and/or flag results on each text instance across all datasets. To get a sense of the general suppression rate of these APIs, we first calculate the false positive rate (FPR) for each API that provides a resulting flag. We find that OpenAI and Llama Guard have similar performance, with FPRs of $0.19$ and $0.20$, respectively, while Anthropic’s FPR is $0.30$.
Looking at the resulting maximum speech suppression scores across identities (summary results in Table 4, per dataset results in Tables 11 and 12) we find that all APIs suffer from identity-related speech suppression for at least one identity group, with some incorrectly flagging identity-related text at 2 or 3 times the overall rate.
@@ 5.2 Speech Suppression results
Figure 1: Identity-related speech suppression values for content moderation API results for flags (top row) and score values (bottom row) for both traditional and generative AI datasets. Error bars are the 95% confidence interval calculated with $1000$ bootstrap samples drawn to the full dataset sizes. Values greater than $1.0$ indicate that text related to that identity is flagged incorrectly more often (top row) or receives higher scores than usual (bottom row) for that dataset and API.
@@
Considering the trends across all identity groups and APIs (see Figure 1 and Tables 13 and 14), we additionally find that while most identity groups suffer from speech suppression under some dataset and content moderation system, there is consistently high speech suppression for speech related to non-Christian religions across APIs. If we compare identity groups within categories (race, gender, and so on), we find that on the traditional content moderation datasets the marginalized group has generally worse speech suppression (except for white versus non-white) across APIs, while the trend is less clear for the generative AI related data.
We also find that the differences between the generative AI and traditional datasets matter for identity-related speech suppression. Some content moderation systems generally have less identity-related speech suppression on some identity groups in the traditional data while others have less suppression for some identity groups on the generative AI data. For example, Google suffers from more non-Christian speech suppression on the generative AI data than the traditional data. Surprisingly, this does not directly align with the goals of these systems; Jigsaw’s Perspective API was designed for toxicity-detection for a traditional content moderation task, yet suffers from more identity-related speech suppression for the LGBT group on the traditional data.
5.3 Regression results
@@
We evaluate the identity-related speech suppression of each API by focusing on cases where they suggest censoring content that should not actually be censored. For the three APIs that return binary flags (OpenAI, Anthropic, and OctoAI) we train logistic regressions to predict whether the flag on a given piece of text matches the ground truth value. For the three APIs that return continuous scores between 0 and 1 (OpenAI—which outputs flags and scores—Perspective, and Google), we train linear regressions that predict higher scores, restricting the data to true negatives—content that, according to ground truth, is appropriate and should not be censored, and where higher assigned scores are more incorrect. In all six models, inputs include our nine identity tags (binary values indicating whether the content pertains to attributes like disability, LGBT, Christians, or women), as well as three other descriptive attributes: whether the text was identity-tagged because it contains a slur or slang term, whether it comes from a generative AI-focused dataset, and its word count.
@@
Across all six models (full statistical results are in Appendix Tables 15 and 16) the immediately apparent result is that most identity-related content is statistically significantly more likely to be incorrectly flagged or incorrectly highly rated than non-identity-related content. This is true for most models, whether the identity tag is for a marginalized group (women, non-Christian, LGBT, non-white, and disability) or for most dominant groups (men, white). Breaking from this trend, however, are a couple notable exceptions: most of the models do not display significant speech suppression towards Christian and straight-related content.
The second major finding from these models pertains to the additional metadata we included in the model. In particular, we found that generative AI-focused text was less likely to be speech suppressed. In the flag-output models, it was more likely to be accurately labeled by Anthropic, OctoAI, and OpenAI, and in the score-output models non-violating content (correctly) received lower scores from Perspective, Google, and OpenAI, though the effect sizes in the latter case were very small. We included the length in words of each instance as a variable, to account for the possibility that differences between generative AI and traditional datasets were due to the generative AI data having a larger average word length. However the effect sizes for word length were very small, indicating that this is not the sole difference between the content moderation of generative AI content versus the traditional task.
6 Discussion and Conclusion
In this paper, we present the first comprehensive audit of automated content moderation systems with a focus on identity-related speech suppression and generative AI. It includes data from more than $570,000$ instances across seven benchmark datasets, including two new datasets, with results from five content moderation APIs and with introduced automatic categorization of this data relating to nine identity groups.

### v3

@@ 5.1. All APIs suffer from identity-related speech suppression
Identity Supp. Identity Supp.
Google non-chr. 1.20 non-chr. 3.58
Jigsaw lgbt 1.58 men 1.22
OAI score non-chr. 3.26 women 1.88
OAI flag non-chr. 2.19 men 5.05
Llama white 2.33 non-chr. 2.35
Anthropic white 1.97 men 2.96
@@
Table 2. Worst identity-related speech suppression values (Supp.) and associated identity group achieving that value (Identity) across content moderation APIs tested. OpenAI (OAI) has speech suppression values associated with both scores (based on $i$-medians) and flags (based on $i$-FPRs) shown, while Google and Jigsaw are both score-based and Llama Guard (Llama) and Anthropic are flag-based.
Speech suppression scores were calculated per-API across all identities and datasets. Summary results showing the identity group and associated worst speech suppression scores per-API are given in Table 2 (with per dataset results in Appx. Tables 12 and 13). We find that all APIs suffer from identity-related speech suppression for at least one identity group, with most incorrectly flagging at least one identity’s related text at 2 or 3 times the overall rate (or with median scores 2 to 3 times as high as the overall median for true negative instances). One API – Jigsaw – however, performs better, with a worst-case speech suppression score of $1.58$ for LGBT-related content on the traditional data and with most identity group scores close to $1$, i.e. parity. This is promising, since Jigsaw has a history of attending to this issue (borkan2019nuanced, ; dixon2018measuring, ; jigsaw-kaggle-unintended-bias, ), demonstrating that steps can be taken by other APIs to reduce identity-related speech suppression. While speech suppression scores are useful for comparing across APIs, it’s useful to understand these scores in the context of the underlying false positive rates; $0.2$ and $0.03$ for OpenAI, $0.2$ and $0.06$ for Llama Guard, and $0.6$ and $0.08$ for Anthropic on the traditional and generative AI data, respectively. For Anthropic this results in an $i$-FPR for men on the generative AI data of $0.247$; in other words, about 25% of content about men in the generative AI dataset that has a label indicating that it should not be flagged as violating is filtered out.
@@
Figure 2. Identity-related speech suppression values for content moderation API results for flags (top row) and score values (bottom row) for both traditional and generative AI datasets. Error bars are the 95% confidence interval calculated with $1000$ bootstrap samples drawn to the full dataset sizes. Values greater than $1.0$ indicate that text related to that identity is flagged incorrectly more often (top row) or receives higher scores than usual (bottom row) for that dataset and API.
[…]
[Identity-related speech suppression results overview]Speech suppression results are shown for six APIs and across nine identity groups separated by the results on the traditional versus generative AI datasets. The worst scores are: men and non-Christian generative AI content for OpenAI’s flags; non-Christian generative AI content and white and non-white traditional content for Llama Guard; men and non-Christian generative AI content for Anthropic; white and non-Christian traditional content for OpenAI’s scores; non-Christian generative AI content for Google; and white and LGBT traditional content for Jigsaw.
@@
Considering the trends across all identity groups and APIs (see Fig. 2 and Appx. Tables 14 and 15), we find that while most identity groups suffer from speech suppression under some dataset and content moderation system, there is consistently high speech suppression related to non-Christian religions across APIs. Additionally, many APIs suffer from speech suppression on both dominant and marginalized groups. For example, Llama Guard has high speech suppression rates for both white and non-white content on the traditional dataset. While there are variations by specific dataset within the traditional and generative AI categories (Appx. Tables 12 and 13), the overall trends hold.
5.1.1. Regression analysis
@@
To further validate the key finding that these APIs suppress identity-related speech, we conducted a regression analysis. Consistent with our speech suppression definitions and approach, we evaluate each API by focusing on cases where they suggest censoring content that should not actually be censored. For the three APIs that return binary flags (OpenAI, Anthropic, and Llama Guard) we train logistic regressions to predict whether the flag on a given true negative labeled text instance results in a false positive (i.e., is incorrectly flagged). For the three APIs that return continuous scores between 0 and 1 (OpenAI, Jigsaw, and Google), we train linear regressions that predict higher scores, restricting the data to true negatives—content that, according to ground truth, is appropriate and should not be censored, and where higher assigned scores are more incorrect. In all six models, inputs include our nine identity tags (binary values indicating whether the content pertains to attributes like disability, LGBT, Christians, or women), as well as three other descriptive attributes: whether the text was identity-tagged because it contains a slur or slang term, whether it comes from a generative AI-focused dataset, and its word count (centered at 0 and scaled to unit variance).
Variable Anthropic Llama Guard OpenAI flag OpenAI score Google Jigsaw
men 0.2837 *** 0.3508 *** 0.4636 *** 0.1642 *** -0.0077 *** 0.0507 ***
women 0.6462 *** 0.7647 *** 0.6872 *** 0.2508 *** 0.0382 *** -0.0861 ***
white 0.6089 *** 1.1958 *** 0.4398 *** 0.1367 *** -0.0908 *** 0.1396 ***
non-white 0.1410 *** 0.8705 *** 0.3231 *** 0.0956 *** -$0 **$1918 ***
christian -0.1488 *** -0.2610 *** 0.2231 *** 0.0718 *** 0.3053 *** 0.0084 ***
non-christian 1.3272 *** 1.6907 *** 1.1227 *** 0.4624 *** 0.0218 *** 0.0844 ***
straight -0.5176 *** 0.0012 -0.0087 -0.0247 ** -0.0893 *** 0.0371 ***
lgbt 0.0085 0.6876 *** 0.5107 *** 0.1609 *** -0.0305 *** 0.1388 ***
disability 0.0545 ** -0.1798 *** -0.1525 *** -0.0416 *** 0.1590 *** 0.0793 ***
GenAI -2.2252 *** -0.8198 *** -3.0622 *** -0.3921 ** -0.4043 *** -0.0625 ***
has_slur -0.6203 *** -0.6222 *** -0.0037 0.0165 *** 0.0494 *** -0.0362 ***
word_length 0.1863 *** -0.5018*** 0.2069 *** 0.0511 *** 0.0442 *** 0.0061 ***
Table 3. Logistic regression results for Anthropic, Llama Guard, and OpenAI flag outputs (APIs outputting binary flags) predicting whether the API gives a false positive: flags on most identity tags are significantly more likely to be incorrectly flagged. Linear regression results for OpenAI score outputs, Google, and Jigsaw (APIs outputting continuous scores) on true negative text (content that should not be moderated) reflect greater speech suppression for identity-related text than other text: most identity tags are significantly more likely to receive higher (worse) content moderation scores. ** $p<0.05$, *** $p<0.005$.
@@
Across all regression models (see Table 3) the immediately apparent result, confirming the above analysis, is that identity-related content is statistically significantly likely to be incorrectly flagged or incorrectly scored higher than non-identity-related text. This is true for most APIs, whether the identity tag is for a marginalized group (women, non-Christian, LGBT, non-white, and disability) or for most dominant groups (men, white). Breaking from this trend, however, most of the APIs do not display significant speech suppression towards straight content, and some are even less likely to suppress straight-related speech than other speech. All other identity groups have one or two APIs which were not found to suppress related speech, with the exception of non-Christian content; consistent with the literature on the difficulty in removing violent associations with Muslim content (abid2021persistent, ), non-Christian content was found to be speech suppressed across all APIs, including some with notably high effect sizes.
5.2. There are differences in identity-related speech suppression for traditional and generative AI datasets
We find that the differences between the generative AI and traditional datasets matter for identity-related speech suppression. Some content moderation systems generally have less identity-related speech suppression on some identity groups in the traditional data while others have less suppression for some identity groups on the generative AI data. For example, Google suffers from more non-Christian speech suppression on the generative AI data than the traditional data (see Fig. 2). Surprisingly, this does not directly align with the goals of these systems; Jigsaw’s Perspective API was designed for toxicity-detection for a traditional content moderation task, yet suffers from more identity-related speech suppression for the LGBT group on the traditional data (although these differences are small given that this model does comparatively well overall). If we compare identity groups within demographic categories (race, gender, and so on, see Fig. 2 and Appendix Tables 14 and 15), we find that on the traditional content moderation datasets, with the exception of race, the marginalized group has worse speech suppression across APIs. The trend is less clear for the generative AI related data. For example, across all APIs on the generative AI dataset white and non-white groups experience similar levels of speech suppression as each other, and the trend is mixed for sexual orientation.
5.2.1. Regression analysis
In order to further test this finding, we included three key variables as part of the regression analysis (described in Section 5.1.1): a Boolean flag indicating whether the instance was text from the traditional or generative AI dataset, a flag indicating whether the instance was identity tagged based on containing a slur or slang term, and a (scaled) count of number of words per instance. Increased word length was found to make almost all APIs (with the exception of Llama Guard) more likely to incorrectly flag. Using the generative AI dataset Boolean flag, we found that generative AI-focused text was less likely to be incorrectly speech suppressed, with large effect sizes, despite that the average word count of the instances in the generative AI datasets is larger. One possible interpretation is that despite the difficulty of correctly flagging longer text instances, the specific datasets contained in the generative AI set were easier content moderation problems. This may especially have been true for the TV and Movies datasets; given that the content comes from Wikipedia, IMDB, and TMDB, we expect that none of it would be particularly inappropriate, and they may contain fewer hard-to-classify instances than the traditional datasets.
5.3. Incorrectly suppressed text is often political speech or TV violence
Traditional Generative AI
API Pol. Rel. Iden. Sex. Hate Viol. None Pol. Rel Iden. Sex. Hate Viol. None
Jigsaw 6 2 12 12 0 2 23 3 2 4 11 0 24 19
Anthropic 35 11 9 3 1 3 8 10 3 3 13 0 32 8
OpenAI 42 17 23 3 3 8 2 10 7 7 25 1 29 1
Llama Guard 26 16 18 5 1 4 14 2 2 2 3 0 11 34
Google 25 5 1 0 1 14 20 5 4 4 6 2 30 15
Table 4. The qualitative categorization of 50 sampled instances per API for both the traditional and generative AI datasets based on categories: politics (Pol.), religion (Rel.), identity bias (Iden.), sexual content (Sex.), hate, violence (Viol.), and no category (none).
To better understand the themes underpinning false positive responses from APIs, we conduct a qualitative review of these responses. We again focus on false positives (for the APIs that return flags) or incorrectly high scoring (for the APIs that return scores) responses. We sample 50 false positive or high scoring texts from each of the generative AI and traditional datasets for each API. For the Google and Jigsaw models, we take our sample from the 1500 highest flagging true negative instances. For each sampled instance, we hand code the instance into zero, one, or more of six categories. Since many models share categorization aims (see Fig. 1) regarding sexual content, violence, and targeted hate, we include these categories in our analysis, and add three additional groups based on our observations of prevalence in a pilot test: religion, politics, and identity bias. For religion and politics, we consider any mention of religious or political figures, systems, or ideologies to count towards their respective categories. We define a text as containing identity bias if it exhibits any biased speech against a named identity group, such as Muslim, LGBT, or Mexican people. Finally, we consider targeted hate to include mentions of violence, threats, or hate speech against specific identity groups.
Based on our qualitative coding, we find different trends for false positive results within generative AI and traditional datasets. On the traditional dataset, Anthropic, OpenAI, and Llama Guard all predominantly flag text regarding politics, religion, and identity bias, the content of which is often argumentative (see examples in Appx. Tables 1618). Google also tends to flag political speech, though this is followed by high flagging rates for violence, not religious or identity biased speech. Jigsaw diverges from these trends in the traditional dataset by flagging uncategorized, identity bias, and sexual content the most; the model largely flags text which uses inappropriate language, even if not as an insult (see Appx. Table 16). For the generative dataset, we find that violence is the dominant category across APIs for false positive text, followed by sexual or uncategorized content. Many of these violent texts are from TV show or movie synopses labeled as PG-13-ok exhibiting violent themes such as shootings, action sequences, or paranormal activity (this finding matches that of previous work on OpenAI in the context of TV shows (mahomed2024auditing, )). Notably, Llama Guard predominantly flags content which does not fall under a qualitative assessment category. Upon reading these texts, we find that the model is often flagging TV and movie descriptions which appear largely benign (see Appendix Table 18).
5.4. Identity groups are incorrectly flagged based on stereotypes and text associations
Figure 3. Comparison of speech suppression scores across different APIs and identity categories (disability, sexual orientation, and religion). Error bars represent 95% confidence intervals based on 1000 bootstrap samples. The religion chart includes a zoomed in version of two of the APIs for clearer comparison.
[…]
[Speech suppression scores by identity and category]Speech suppression scores for disability, sexual orientation, and religion are shown for APIs that have category-specific scores or flags (OpenAI, Llama Guard, and Google) across commonly shared categories associated with identity group stereotypes or text associations (disability is analyzed based on self harm, sexual orientation based on sexual content, and religion based on hate or toxic speech).
In order to identify why identity-related speech suffers from suppression on both the traditional and generative AI datasets, we further consider the speech suppression scores per-identity and per-category across the traditional and generative AI datasets (see Appx. Figures 710). We find that text pertaining to different identity groups are flagged for different reasons. For example, when the Google API is applied to the generative AI dataset, disability-related content receives a high speech suppression score and is flagged for “Health” and non-Christian content is flagged for “Religion & Belief.” On the Llama Guard API we find that generative AI content related to men is incorrectly flagged for “Violence & Hate” and for “Sexual Content.” From the literature (see Sec. 2) we know that different identity groups have associations in most large language models, often based on stereotypes, that must be purposefully unbiased if not desired.
To examine these further, we consider the speech suppression scores for the identity categories of disability, sexual orientation, and religion across APIs that return per-category scores or flags (OpenAI, Llama, and Google). We consider categories that are relatively consistent in flagging aim across API and for which we anticipate a stereotypical or other textual association with the identities based on the literature. In Figure 3 we see that across APIs: disability-related content is more likely to be flagged for self-harm than other content; straight and LGBT content is more likely to be flagged for sexual content than non-identity-related content, and; non-Christian content is more likely to be flagged for hate than Christian or non-identity-related content.
These results are somewhat expected based on known stereotypes, but also provided needed specificity to the literature. LGBT content is stereotypically associated with inappropriate sexual content, but we find here that both straight and LGBT content are incorrectly flagged by these systems as sexual. It’s known that Muslim content is commonly associated with violence by large language models (abid2021persistent, ), but these results additionally provide initial evidence that non-Christian content is heavily associated with hateful speech, leading speech related to these groups to be incorrectly flagged as hateful. While disability content is known to be identified with negative sentiment (venkit2022study, ), examining the per-category flags also allows us to identify the association between disability and speech suppression related to self-harm. Much of the literature on bias focuses on a binary identification of toxic content (see, e.g., (dixon2018measuring, ; borkan2019nuanced, )), and thus may miss some of this per-identity and per-category nuance. Examining these stereotypes and text associations in this fine-grained way may allow future avenues for bias reduction to be identified.
Table 5. Jigsaw Bias Speech Suppression Scores
Identity group Google Jigsaw OpenAI OpenAI Llama Anthropic
(scores) (flags) Guard
men 0.33 0.46 0.56 0.0 0.16 1.91
women 0.28 0.49 0.65 0.0 0.15 1.56
straight 0.93 2.28 2.55 1.83 0.41 0.63
lgbt 1.16 3.6 2.26 3.59 0.89 0.7
non-white 0.42 0.67 0.81 0.0 1.75 0.81
white 0.59 1.29 0.72 0.0 1.15 5.15
christian 2.16 0.37 0.35 0.0 0.27 0.17
non-christian 2.16 0.57 0.39 0.1 0.71 0.85
disability 1.65 2.92 1.8 0.0 0.16 0.89
Additionally, to better understand the extent to which these stereotypes and negative textual associations that lead to incorrect flagging behavior are based on the identity terms themselves or the broader context of the examined text, we turn to the speech suppression results on the Jigsaw Bias dataset (see Table 5). Recall that this dataset is created in a template-based manner, with different explicit identity terms substituted into the same phrases. This allows us to see what the speech suppression scores are like across identity group and API in cases where the broader context of the text has been removed and suppression is likely due to the identity terms themselves. We see that all APIs still suffer from speech suppression for some identity groups on this dataset. Interestingly, OpenAI’s content moderation endpoint was directly trained using this template-like approach (markov2023holistic, ). We see that while this brought the speech suppression for most identity groups on the dataset to essentially 0 for OpenAI’s flag-based results, this was notably unsuccessful for the LGBT and straight groups. Across APIs, this helps us to understand that explicit identity references are associated with some portion of the identified speech suppression, in addition to further suppression based on the text content.
6. Discussion and Limitations
Our focus on identity-based speech suppression by LLM content moderation endpoints represents a novel contribution to the content moderation literature. While most content moderation and AI safety research prioritizes preventing harmful content from passing through filters (false negatives), we argue that examining false positives is also important, particularly in generative AI contexts where creative expression may be impacted.
6.1. Balancing false positives and false negatives
While content moderation endpoints for LLMs are trained for company-specific standards for text toxicity, our TV and movie plots datasets represent normative judgments about content appropriateness derived from established rating systems. While we recognize that companies must establish their own content moderation norms aligned with their values and risk tolerances, we believe it is important to audit how these privately trained content moderation filters might diverge from widespread norms for appropriate content (such as the age ratings we use for the TV and movies dataset). Determining the appropriate balance between company-specific priorities and public expectations of content moderation remains an open question.
We also recognize the legitimate, prosocial reasons companies may “err on the side of caution” by over-filtering text. As described in Section 2.3, generative AI systems have produced content that has harmed and endangered real people. The public focus on harm creates well-justified pressure to prioritize preventing false negatives, even at the cost of increased false positives. Nevertheless, our results suggest that these safety-oriented approaches may have unintended consequences for creative expression, particularly for content related to identity.
6.2. Innovations and limitations
One innovation of this work is the use of “in-the-wild” datasets and associated labels — the TV and movie datasets with corresponding age ratings — that we introduce to assess the impact of automated content moderation on creative content. Our focus on identifying falsely flagged content allows us to make use of this human-created content, providing a benchmark for normative societal expectations in these domains. At the same time, this presents a limitation: the nature of the TV and movies dataset prevents us from attributing sole causality for this over-flagging to the presence of particular identity groups in the text, as there are unaccounted-for differences in the distribution of content across TV and movie synopses containing references to different identity groups. It is exactly these differences we are interested in assessing. We can gain some insight into controlled differences through the Jigsaw Bias dataset, which was artificially constructed by substituting identity words into short phrases. The presence of speech suppression even on this dataset demonstrates that references to identity groups on their own must be at least some source of the over-flagging of identity-related speech.
One site of possible disagreement and limitations of this work stems from the foundational question of what it means for a text instance to be “identity-related.” In this work, we use a number of schemes to tag specific datasets by identity: template-based, in which the identity is explicitly referenced (e.g., “hug gay”); individually coded, in which the text has been manually identified by a person as related to the identity; text references, in which we identify a text as related to an identity if it includes a slur, slang, or neutral term related to that identity; and external associations, in which we find external information associating TV shows and movies with identity groups. We use multiple tagging methods to mitigate the limitations of any one of these methods, casting a wide net to identify “identity-related speech”, and find speech suppression across all tagging methods.
The specific flagging and stereotype patterns we identify in identity-related speech suppression are likely influenced by both the identity tagging schemes and the specific datasets and labels used. For example, the regression results (see Section 5.2.1) indicate that the automated content moderation systems tested are more poorly calibrated in flagging longer content, yet these systems more accurately flag the longer generative AI dataset, likely because it contains content that is more “fluent” and easy to categorize relative to the shorter user-generated comments that make up much of the traditional dataset. When considering the performance of APIs on specific identities, it is similarly hard to disentangle the co-occurrence of identity-related and harmful content. The analyses in Sections 5.3 and 5.4 assess and disentangle some of these effects, and some are mitigated through the focus throughout on incorrect suppression, but there may be additional co-occurrence effects that future work could usefully identify.
6.3. Future work
Achieving both precise identification of unwanted content and identity-based fairness in alignment with any standard is an unsolved technical challenge. After all, human-facilitated content moderation has well-documented difficulties with efficacy and fairness across identity groups (see Section 2); we should therefore expect that, even with good-faith efforts by companies to address fairness issues, fully automated content moderation filters will struggle with both efficacy and fairness.
We see two possible useful paths forward. First, non-automated content moderation has long employed humans to help moderate content in cases of ambiguous policy application, with concerning job conditions and mental health outcomes (gray2019ghost, ; roberts2019behind, ). While finding ways to incorporate humans-in-the-loop to screen generated items may therefore be not only impractical for generative AI but also inadvisable, there may be ways to incorporate the desires and opinions of generative AI users via carefully designed preferences or interface interactions. Future work could consider how best to draw on user input as part of this process.
Second, with the identification that identity-related text is flagged based on stereotypes and text associations that are specific to different demographic groups (see Section 5.4), there may be targeted approaches that can be taken to decouple incorrectly flagged identity-related text from notions of undesired content. For example, we saw that OpenAI had some success by using a template-based explicit identity term approach in training to overcome some (but not all) undesired text associations when tested on a similar template-based dataset. Purposeful training to remove such associations and stereotypes from these automated content moderation systems may be effective in increasing both accuracy on these incorrectly flagged instances and fairness in the distributions of such errors.
Finally, we also encourage future work that further examines the impact of identity-related speech suppression in creative contexts, for example by working directly with creators. Such work could also usefully examine how these APIs are built on by other apps or software layers, and how these choices impact the overall user experience and resulting creative work.
7. Conclusion
Our research shifts focus from the traditional concern of content moderation — filtering out toxic or undesired content — to examine the rate and patterns of false positives: non-toxic content incorrectly flagged by moderation systems. While it is crucial to filter harmful content, the incorrect suppression of legitimate speech carries significant consequences, particularly when unevenly distributed across identity groups. We introduce speech suppression scores to quantify this disparity, measuring the ratio between the false positive rate for text referencing particular identity groups versus the overall false positive rate.
To enable a comprehensive analysis of speech suppression in contexts relevant to generative AI, we developed a novel dataset of TV and movie plots with their associated content ratings as a naturalistic measure of content appropriateness. This dataset provides several advantages over traditional content moderation datasets like Jigsaw and TweetEval. While traditional datasets focus on short-form user-generated content from social media and comment sections, our TV and movie plots represent longer-form creative content of the type increasingly produced through generative AI systems.